OpenStudy (anonymous):

Morissa pays $11.29 per pound for dried blueberries and $10.19 per pound for dried cherries. Which of the following statements could be true?

OpenStudy (anonymous):

Which statements ?!:)

OpenStudy (anonymous):

a.Morissa bought 1.3 pounds of blueberries and 1.7 pounds of cherries for a total of $32. b. Morissa spent $22.58 on blueberries and spent $20.38 on cherries for a total of 3 pounds of fruit. c. Morissa bought 2 pounds of fruit for a total of $19.98. d.Morissa bought 2 pounds of blueberries and spent $9.99 on cherries for a total of $34.99.

OpenStudy (anonymous):

He spent for : a pound of blueberries =2*10/19 / / / / cherries =1/7

OpenStudy (anonymous):

The answer of a statement that is true should be A

OpenStudy (anonymous):

Thanks Sooo Much!!!! ;)

OpenStudy (anonymous):

To both of you!!

OpenStudy (anonymous):

Yeah ! A is true !:) :D